OpenStudy (anonymous):

find degree and radian of function value of sec = sqrt(2)

OpenStudy (anonymous):

help pls

OpenStudy (anonymous):

rad = 6.412570906 deg=1.000304695

OpenStudy (anonymous):

So you're looking for the angle whose secant is \[\sqrt{2}\]. \[\sec \theta\]\[ = 1/\cos (\theta)\], so the reciprocal is true, too. \[\cos \theta=\frac{ 1 }{ \sec (\theta) }\] \[\cos \theta =\frac{ 1 }{ \sqrt{2}}\] Where on the unit circle is cosine \[\frac{ 1 }{ \sqrt{2} } or \frac{ \sqrt{2} }{2}\]? \[\frac{ \pi }{ 4}\] radians or 45 degrees